A scientist claims that only 65% of geese in his area fly south for the winter. He tags 55 random geese in the summer and finds that 17 of them do not fly south in the winter. If = .10, is the scientist's belief warranted?

A. Yes, because the test value 0.64 is in the noncritical region.

B. No, because the test value - 5.30 is in the critical region.

C. Yes, because the test value –0.30 is in the noncritical region.

D. No, because the test value –0.64 is in the noncritical region.

Answer by stanbon(75887)

A scientist claims that only 65% of geese in his area fly south for the winter. He tags 55 random geese in the summer and finds that 17 of them do not fly south in the winter. If = .10, is the scientist's belief warranted?
Ho: p <= 0.65 (claim)
Ha: p > 0.65
sample proportion = 28/55 = 0.69
z(0.31) = (0.69-0.65)/sqrt(0.65*0.35/55) = -5.29
p-value = p(z < -5.29) = normalcdf(-100,-5.29) = 0.64
Since the p-value is greater than 0.10, fail to reject the claim
A. Yes, because the test value 0.64 is in the noncritical region.
-----
Cheers,
Stan H.
-----------